### v2

# Breast Histopathology Image Retrieval by Attention-based Adversarially Regularized Variational Graph Autoencoder with Contrastive Learning-Based Feature Extraction

Abstract
Breast cancer is the most common cancer type in women worldwide. Early detection and appropriate treatment can significantly reduce its impact. While histopathology examinations play a vital role in rapid and accurate diagnosis, they often require a substantial workforce and experienced medical experts for proper recognition and cancer grading. Automated image retrieval systems have the potential to assist pathologists in identifying cancerous tissues, thereby accelerating the diagnostic process. Nevertheless, due to considerable variability among the tissue and cell patterns in histological images, proposing an accurate image retrieval model is very challenging.
This work introduces a novel attention-based adversarially regularized variational graph autoencoder model for breast histological image retrieval. Additionally, we incorporated cluster-guided contrastive learning as the graph feature extractor to boost the retrieval performance. We evaluated the performance of the proposed model on two publicly available datasets of breast cancer histological images and achieved superior or very competitive retrieval performance, with average mAP scores of 96.5% for the BreakHis dataset and 94.7% for the BACH dataset, and mVP scores of 91.9% and 91.3%, respectively.
Our proposed retrieval model has the potential to be used in clinical settings to enhance diagnostic performance and ultimately benefit patients.
keywords:
Computational pathology , Graph Neural Network , Graph Construction , Adversarial Training , Variational Graph Autoencoder , Attention Module , Image Retrieval
1 Introduction
Breast cancer is a global health concern for women and has led to considerable mortality worldwide [1]. It is the most commonly diagnosed cancer and represents around 25% of all cancers diagnoses in women [2]. According to 2018 statistics, breast cancer accounted for 15% of all cancer-related deaths and has increased by 24% in recent years [3]. While breast cancer causes many deaths worldwide, early detection and diagnosis followed by appropriate treatment can help to reduce its impact [4]. The need for early detection of breast cancer has led to new solutions, such as computer-aided diagnosis systems based on histological images [5]. This technology is a valuable tool to improve breast cancer diagnosis and potentially can save lives [6]. However, proper diagnosis based on histological images often requires a substantial workforce and experienced medical experts. To overcome this challenge, pathologists can benefit from referring to archived databases for complex cases to find similar examples with verified findings to minimize diagnostic errors.
Content-based image retrieval (CBIR) is an approach to identify visually similar images in a database and can be used for various applications [7, 8]. For example, pathologists reviewing tissue sections for breast cancer diagnosis may benefit from searching through similar labeled histopathological image data to improve their diagnostic accuracy in distinguishing cancerous from non-cancerous tissue. CBIR techniques have also been used for various non-medical applications such as ”image search” for general and natural images, visual search for retail products, and multiple tools for recognizing faces and artworks [9]. Prior machine learning-based CBIR systems often employed models adapted to specific applications, requiring the collection of labeled data for each particular use, which presented a significant implementation challenge. Additionally, the term ”similarity” in these approaches could have different interpretations depending on the specific applications [10]. For instance, two histological images might be considered similar because they originate from the same organ, are classified as the same type of cancer, share similar staining, or represent comparable histological features.
Content-based medical image retrieval (CBMIR), a subset of CBIR, is specifically adapted to medical images and has been used in the context of radiology and histology in former studies [9, 10, 11]. CBMIR for histological images, not only aids histopathology in grading new tissue samples but also assists in analyzing patterns within similar tissues from previous patients [11, 5]. CBMIR offers distinct advantages over classification tasks. For instance, in classifying prostate cancer, a classification model provides class labels for each input. In contrast, a CBMIR system produces the top-$k$ most similar images based on a query image, where $k$ represents the number of desired results. Upon availability, it can also provide other relevant information such as clinical meta-data or diagnosis reports [12]. Therefore, CBMIR provides pathologists and practitioners with more comprehensive information about diverse tissue patterns and enables them to identify the most similar patterns within different tissues, which could potentially minimize intra-and inter-observer variability [9, 13]. Another advantage of CBMIR is that it inherently includes a clinician-in-the-loop, which is often required for artificial intelligence (AI) applications in healthcare. Furthermore, a pathologist using CBMIR in the diagnostic process is required to come to their own diagnostic conclusion after incorporating additional information provided by CBMIR systems rather than replacing the pathologist’s diagnosis with an AI-based classifier. Unlike histological images with their reports written by pathologists that are presented by CBMIR systems, a classifier is not necessarily easily interpretable and - even if presented to a clinician as a second opinion - may lead to automation bias [14, 15]. From a machine learning perspective, CBMIR is typically performed using unsupervised techniques, while classifiers are usually trained with supervised methods that require labeled datasets.
Histopathology image retrieval is a very challenging task due to the substantial tissue and cell pattern variability [10]. Various semi- and fully-automatic approaches have been proposed for histopathology image retrieval [9]. Among these, graph neural networks (GNN) have gained considerable attention in recent years. Considering the grid data structure in images and the tremendous developments in convolutional neural networks (CNNs), many graph convolutional neural network (GCNN)-based architectures have also been extensively proposed for a wide range of computer vision tasks, including image retrieval [16]. In GCNN, input images can be represented as the graph nodes, and their features are usually extracted by pre-trained CNNs such as Visual Geometry Group (VGG) [17] or EfficientNet [18] to construct the graph. However, the exploited pre-trained models are typically trained on natural images, and their extracted features may not be optimally suited for medical data (e.g., histological images) [19, 20].
While supervised learning is usually used to train GCNNs, some GCNN variants, such as variational graph autoencoder (VGAE) [21], can be used for unsupervised learning tasks. In VGAE, the encoder network maps input data to a lower-dimensional latent space, and the decoder network maps latent space vectors back to the original input data. Unlike conventional graph autoencoders (GAE), VGAE utilizes probabilistic encoding that maps input data to a probability distribution over the latent space. VGAE performance can be further improved by incorporating other techniques, such as adversarial training or adding an attention mechanism to the model [21]. In adversarially regularized variational graph autoencoder (ARVGA) [22], the encoder is forced to generate embedding vectors resembling the data from a prior distribution through adversarial training. Incorporating an attention mechanism into CNNs [23] or GNNs [24] is another technique that has been exploited. These techniques have also been jointly applied for specific applications such as graph link prediction [25], graph clustering visualization [25], and node classification in partially labeled graphs [26].
In this study, for the first time, we adapted and exploited attention-based adversarially regularized variational graph autoencoders (A-ARVGAE) for the histological breast image retrieval task. Moreover, unlike previous studies that use conventional pre-trained CNNs for feature extraction to build the graph, we used the recently developed clustering-guided contrastive learning (CCL) feature extractor [27], which is explicitly designed for histopathological images. We evaluated the performance of our proposed approach on two binary and multi-class histopathological breast image datasets, namely breast cancer histopathological database (BreakHis) [28] and breast cancer histology images dataset (BACH [29]), and achieved superior or very competitive performance compared to other state-of-the-art models. The main contribution of this research can be summarized as follows:
  • 1.
    We adapted and exploited one of the state-of-the-art GNN models, which incorporated both attention mechanism and adversarial training for breast cancer histology image retrieval task.
  • 2.
    Instead of using conventional pre-trained CNNs as the node feature extractor, we made use of CLL, which is explicitly designed for histological images.
  • 3.
    Through extensive experiments, we show the excellent performance of our approach for both utilized GNN and feature extraction on two histopathological breast image datasets, namely BreakHis and BACH, and achieved superior or very competitive performance compared to other approaches.
The rest of this paper is organized as follows: In Section 2, we present a background on methods for medical image retrieval and graph representation learning in histopathology. Section 3 describes our methodology, which contains all steps for the proposed framework. Section 4 consists of the experiments and evaluations with an analysis of the results. Finally, in Section 5, we conclude the paper with suggestions for future works.
2 Related Works
This section contains a literature review for medical image retrieval and graph representation learning in histopathology, considering their relevance to our work.
2.1 Medical Image Retrieval
Various methods have been proposed in the literature for medical image retrieval. Text-based image retrieval systems were among the first attempts to find similar images in a given database. However, these methods rely heavily on expert annotation and proper keyword selection, which makes them generally considered ineffective and time-consuming [30].
CBMIR models, on the other hand, rely on the content (features) of the images instead of their given labels. Consequently, they can be used in an unsupervised manner without the need for manual labeling. Conventional CBMIR models consist of image preprocessing, hand-crafted or automatic feature extraction, and feature similarity measurements using different distance-based methods. To enhance interpretability, some approaches propose reducing the disparity between the extracted features and the high-level semantic meanings of the images by incorporating expert feedback or medical metadata into the retrieval model [30, 31].
Commonly used preprocessing methods include noise reduction, contrast enhancement, patch generation (e.g., from whole slide images), and normalization (e.g., stain normalization in histological images) [32, 33]. However, many novel machine learning (ML)- and deep learning (DL)-based feature extractors do not require any preprocessing steps.
For feature extraction, both manually designed and DL-based automatic feature extractors have been utilized in former studies. Hand-crafted features are employed to depict the morphological and textural characteristics of images. These features include color histogram-based features, histogram of oriented gradients (HOG) descriptor, local statistical features, global image structure tensor (GIST) features, Gabor features, wavelet-based features, intensity-based features, textural features, gray level co-occurrence matrix (GLCM) features, and scale-invariant feature transform (SIFT) descriptors [30, 27]. It should be noted that experts usually need to set up specific transformation functions to derive these features. However, more recently, the rapidly evolving field of DL has made it easier to automatically extract high-level features from the images [10].
Many pre-trained DL-based models, such as VGG19, ResNet, MobileNetV2, Xception, EfficientNet, and DenseNet121, have been used as feature extractors for various computer vision tasks such as medical image classification or medical image retrieval [34, 35]. Hand-crafted or automatically extracted features of a query image are then compared to the image features in a database to find similar images. For this comparison, various similarity measurements such as distance-based retrieval (SDR), customized query approach (CQA), or nearest-neighbor matching have been used [36]. In the context of histopathology, Yottixel [37] and FISH [13] methods select the most distinctive patches to represent whole slide images (WSIs) through $K$-means clustering based on RGB histogram and spatial coordinate features. Yottixel employs a pre-trained DenseNet121, trained on the ImageNet dataset [38], to extract patch embeddings, which are then transformed into binary barcodes. FISH trains a model based on an autoencoder on the cancer genome atlas (TCGA) dataset to encode an integer index for each patch during retrieval. However, it is worth mentioning that using a feature extractor trained only on natural images might be sub-optimal due to the fundamental differences between natural and histopathological images. In contrast, CCL is a newly developed model specifically designed for histological images. CCL pretraining was performed with around 15 million image patches extracted from more than 32,000 WSIs.
The CCL learning method utilizes large-scale unlabeled histopathological images for universal feature extraction that can be directly applied for various WSI analysis tasks such as retrieval. In our study, we utilized CCL as a robust and relevant feature extractor to build the graph.
2.2 Graph Representation Learning in Histopathology
While CNNs have demonstrated impressive performance in medical image analysis, they often have trouble at capturing contextual information from neighboring areas, especially with a limited field of view. In contrast to CNNs, GNNs are better at preserving complex correlations among neighboring elements during the learning process. GNNs have been successfully applied to various tasks, including tumor/object localization, classification, survival prediction, and image retrieval [39].
In the field of computational pathology, several GNN models, such as cell graphs, tissue graphs, cell-tissue graphs, and patch graphs, have been proposed for tasks like node classification (e.g., nuclei classification) or graph classification (e.g., histologic image patch categorization) [40, 41]. However, these applications have primarily been applied in a supervised manner and often need advanced preprocessing. Using other models like HoVerNet [42] or DDU-Net [43] to generate nuclei segmentation masks as input for the graph-based model is an example of such preprocessing. Regarding clinical applications, GNNs have been used in different cancer types analysis, including breast cancer, colorectal cancer, and prostate cancer, among others [41].
To build a graph, one needs to define nodes, node embeddings, and edges. In the context of histological images, nodes could represent nuclei, image patches, or tissue regions. Node embeddings are features of the nodes that can be extracted either manually or automatically using pre-trained models or other techniques. The edges depict connections between nodes, typically determined by predefined criteria such as $k$-nearest neighbors or distances between nodes. Once the graph is built, various GNN or GCN models can be applied, depending on the specific applications.
A GCN contains several key components, including an aggregation layer to define how features from neighbor nodes are aggregated, an activation layer to incorporate non-linearity into the model, graph pooling to reduce graph size (if needed), and graph convolutional operations to update node representations. It’s worth noting that other variants, such as GraphSAGE [44], GAT [45], or ChebNet [46] have been introduced in the literature as novel propagation rules to enhance performance across different applications. The output of a GNN typically depends on the application. For graph classification, a multilayer perceptron (MLP) can be added to the graph. In the case of graph reconstruction for unsupervised tasks, a linear projection decoder can be added to the model. As highlighted in a recent review [39], most GNN-based works in computational pathology are related to supervised classification tasks. Only a few studies have investigated histological image retrieval [39, 27, 47, 12]. However, even in these studies, the potential of attention mechanisms and adversarial training have not been explored. In our research, for the first time, we adapted and made use of both attention mechanism and adversarial training in variational graph autoencoders for the histological image retrieval task.
3 Methodology
The general workflow of our proposed approach for histological image retrieval is depicted in Figure 1. In the following, we discuss the different parts of the workflow in detail.
Figure 1: The general workflow of the proposed breast histological image retrieval model involves distinct processes for training and testing. A-ARVGAE: Attention-based adversarially regularized variational graph autoencoders.
3.1 Dataset
The dataset refers to a collection of histological image patches used to train and test the model. Each image has certain features and characteristics related to a specific class or histopathological grade. We use two breast histopathology image datasets in this study, namely the BreakHis dataset [28] and the BACH dataset [29], which are subsequently detailed. Example images from the datasets are shown in Figure 2.
  • 1.
    BreakHis: This dataset contains hematoxylin and eosin (H&E)-stained breast histology microscopy images captured at various magnifications (40×40$\times$40 ×, 100×100$\times$100 ×, 200×200$\times$200 ×, 400×400$\times$400 ×) and is categorized into two primary classes, including benign and malignant classes. The dataset contains 7,909 images including 2,480 benign and 5,429 malignant samples with a size of $700\times 460$ pixels in Portable Network Graphics (PNG) format.
  • 2.
    BACH: This dataset contains H&E-stained breast histology microscopy images captured at 200×200$\times$200 × magnification. It includes 400 images distributed across four distinct classes, namely benign (100 images), in situ carcinoma (100 images), invasive carcinoma (100 images), and normal (100 images), each with a resolution of $2048\times 1536$ pixels in Tagged Image File (TIF) format.
BreakHis(40$\times$) BreakHis(100$\times$) BreakHis(200$\times$) BreakHis(400$\times$) BACH(200$\times$)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Example images from the BreakHis (first four images) and the BACH (last image) datasets.
3.2 Feature Extraction
The first step in our proposed retrieval model is the extraction of features for all images in the dataset. Instead of using hand-crafted features or conventional pre-trained CNNs as feature extractors, we utilize CCL as the main feature extractor. Unlike other conventional pre-trained CNNs, CCL is trained with histological images. We hypothesize that using such a pre-trained feature extractor could lead to better descriptive features for the task of histological image retrieval.
We also use a set of standard baseline models as feature extractors and compare their retrieval performance with the CCL feature extractor. For this, we employ well-known pre-trained models, namely VGG19 [17], MobileNetV2 [48], DenseNet121 [49], Xception [50], InceptionResNetV2 [51], NASNetLarge [52], and EfficientNetV2M [18]. The details of these baseline pre-trained models are presented in Table 1. It is worth mentioning that all these pre-trained models are initially trained on natural images using the ImageNet dataset [38], but the CCL feature extractor used in this study is pre-trained with histological image patches. The results of these comparisons are reported in Section 4 (Table 2 and Table 3).
Table 1: Details of baseline feature extractors. All these pre-trained models are initially trained on natural images using the ImageNet dataset [38].
Feature Extraction Model Size (MB) Parameters Depth
VGG19 549 143.7 M 19
MobileNetV2 14 3.5 M 105
DenseNet121 33 8.1 M 242
Xception 88 22.9 M 81
InceptionResNetV2 215 55.9 M 449
NASNetLarge 343 88.9 M 533
EfficientNetV2M 220 54.4 M -
3.3 Graph Construction
GNNs are mostly applied on graph-based datasets. These datasets typically include nodes, node features, edges, and adjacency matrices. In contrast, non-structural datasets, such as image-based data, lack inherent graph structures. This makes graph construction an essential step for applying GNNs to such data. For the retrieval task, graph construction methods often treat subjects (e.g., images) as graph nodes ($V$), with the features of these subjects acting as node features ($X$). Edges between nodes ($E$) can be defined using distance metrics like cosine and Euclidean distances, indicating the likelihood of interaction between entities [53, 54, 55, 56]. The topology of these graphs is usually determined experimentally, employing strategies such as a pre-defined proximity threshold, a nearest neighbor rule, or a probabilistic model. To create edges between nodes, various algorithms such as $k$ approximate nearest neighbor ($k$-ANN) can be used. This approach is more time-efficient compared to the $k$-NN algorithm and is beneficial, especially in high-dimensional feature spaces, where the complexity of high dimensions slows down exact nearest neighbor searches. While there are various implementations of ANN algorithms, the ANN benchmark [57] is a widely-used library that encompasses most ANN-based methods. In our study, we utilize the fast library for approximate nearest neighbors (FLANN) to calculate distances between node embeddings, selecting the $k$ nearest nodes to each entity as adjacent nodes. The computation of graph edges is conducted as follows:
$e_{ij}=\begin{cases}1&\text{if }v_{j}\in k\text{-ANN}(v_{i})\\ 0&\text{otherwise}\end{cases}\quad(1)$
where $e_{ij}=1$ indicates the existence of an edge connection between the $i$-th and $j$-th nodes in the graph, and $e_{ij}=0$ denotes no edge. Utilizing the edges and nodes, the adjacency matrix ($A$) can be constructed where the values of the matrix ($a_{ij}$) act as indicators, quantifying of the connections between nodes $V_{i}$ and $V_{j}$.
It is worth mentioning that we use transductive learning, where the entire dataset (training, validation, and test data) is observed to build the graph. The summarized workflow to build the graph in our study is shown in Figure 3.
Figure 3: The framework of graph construction. CCL: Cluster guided contrastive learning, $K$-ANN: $K$ approximate nearest neighbors, FLANN: fast library for approximate nearest neighbors
3.4 A-ARVGAE
The overall architecture of the A-ARVGAE model is depicted in Figure 4. The proposed architecture comprises four main components: the encoder (yellow block), which also serves as the generator for adversarial training; the embedding latent space (pink block); the decoder (blue block); and the discriminator (green block), which is another essential component for adversarial training.
The constructed graph, as described in the previous section, is fed into the encoder of the A-ARVGAE model. The encoder includes a graph attention layer that performs attention operations, batch normalization, leaky ReLU, and dropout, followed by two graph convolutional layers. The attention-based convolutional operation integrating attention mechanisms into convolutional layers assigns varying weights to the neighbors, enabling the model to concentrate on the most relevant parts of the graph when learning node representations [45]. The attention mechanism aggregates information from adjacent nodes with an adaptive weight allocation to neighboring nodes. As indicated by prior studies, the use of a multi-head attention layer has been shown to yield superior performance [25] and hence we also employ two-head attention layer in our work.
Inspired by the VGAE [19], the encoder in our architecture does not encode an input into a fixed latent representation. Instead, the outputs of the graph convolutional layers are parameters (mean ($\mu$) and variance ($\sigma$)) that define a probability distribution over the possible latent representations, denoted as $Z$, characterized by a normal distribution $N(\mu,\sigma)$. Sampling from the latent space involves drawing samples from the distribution $N(\mu,\sigma)$ to produce latent node embeddings. This process employs the reparameterization trick, as introduced in [19], which enables the backpropagation of gradients through the stochastic sampling process for gradient descent optimization.
The decoder block employs a linear projection through a straightforward inner product ($Z*Z^{T}$) followed by a sigmoid function, which results in the reconstructed graph.
As the A-ARVGAE model incorporates adversarial training, it includes a generator and a discriminator. Within this framework, the encoder functions analogously to a generator in an adversarial network [22]. The generator tries to deceive the discriminator by producing synthetic data, which, in this context, refers to latent variables generated from the encoding of graph data. On the other hand, the discriminator (green block) attempts to differentiate whether samples are derived from real data or the generator. The discriminator evaluates embeddings from the prior distribution (a Gaussian distribution) as real and those from the latent variable $Z$ as synthetic or fake. The discriminator block comprises a straightforward MLP with two linear layers along with batch normalization, leaky reLU, and dropout to distinguish between real (a prior Gaussian distribution) and fake (latent space $Z$) embeddings.
The complete model integrates three loss components: a regression-based reconstruction loss (related to VGAE), a Kullback–Leibler divergence loss (related to VGAE), and a binary cross-entropy loss (related to discriminator in adversarial training). Further details regarding the model architecture and cost functions can be found in the respective studies [22, 45].
We compare the results of our employed GNN model (A-ARVGAE) with other GNN-based models, namely GAE, VGAE, and ARVGA, and also with another non-GNN-based state-of-the-art approach, called RetCCL, represented in [27]. RetCCL is a model that is recently developed for the histological image retrieval task, which is based on CCL feature extraction and nearest neighbor search to find similar images in a given dataset. The results of these comparisons are reported in Section 4 (Table 4).
Figure 4: Architecture of attention-based adversarially regularized variational graph autoencoder (A-ARVGAE). For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
3.5 Trained Graph Encoder
After training, the trained A-ARVGAE encoder can transform any input image into a vector in the embedding space.
3.6 Image Embedding Database
The embeddings generated by the trained A-ARVGAE encoder in the previous step for each training image are stored in a database. This database is then utilized in the testing phase to identify the top-$K$ similar images for a given query image as explained in Section 3.7
3.7 Testing phase
As depicted in the lower part of Figure 1, for a query image, the model aims to identify the top $K$ similar images in the database. Similar to the process with training images, features are first extracted from the test image using the CCL feature extractor. These features are then fed into the trained A-ARVGAE model to obtain the embedding in a lower-dimensional space, specifically from the encoder part of the model. After acquiring the embedding of the query image, it is compared with the embeddings in the image database to identify similar images. To find the most similar images, the ranking algorithm evaluates and ranks the candidate images based on Euclidean distance. These selected images are then arranged and ranked in ascending order according to Euclidean distances. Once ranked, the top-$K$ similar images are chosen as the final output of the retrieval model.
To evaluate the performance of the developed approach, we use two evaluation metrics, namely, mean Average Precision (mAP) [58] and mean Majority Vote (mMV) [13].
mAP is the average of the precision values computed at the ranks of relevant images in the retrieved results. It is a widely used metric for evaluating the effectiveness of a retrieval system. The formula for mAP is as follows:
$mAP(k)=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\frac{1}{|R_{q}|}\sum_{k=1}^{|R_{q}|}\text{Precision}(k)\cdot\text{rel}(k)\quad(17)$
Where $|Q|$ is the number of queries, $R_{q}$ is the number of relevant images for query image $q$, $\text{Precision}(k)$ is the precision at rank $k$ (the ratio of relevant images retrieved in the top $k$ results to the total number of retrieved images up to rank $k$), $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise. ||$\cdot$|| ⋅ | denotes the length of the set.
mMV evaluates the frequency with which the primary diagnosis among the top-$k$ results aligns with the queried diagnosis as follows:
$mMV(k)=\frac{1}{|Q|}\sum_{q=1}^{|Q|}1\{L_{q}\in MV(\text{rel}(k))\}\quad(18)$
Where $L_{q}$ represents the ground truth diagnosis of the image, and MV is the predicted diagnosis of an image taken from the majority vote of top-$k$ samples. Similar to mAP, $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise. Here, ||$\cdot$|| ⋅ | also denotes the length of the set.
4 Results & Discussion
In this section, we start with the implementation details, then delve into the experimental setup, and finally present the obtained results along with a detailed discussion.
4.1 Implementation Details
We used the PyTorch deep learning framework (version 2.2.0) and the PyTorch Geometric library (version 2.4.0) to implement the models. Adam optimizer was used to train the models with an initial learning rate (LR) of 0.0001. For dropout layers, a probability of 0.2 was chosen in the model architecture. All models were trained for 250 epochs, whereas in models with adversarial components, the discriminator was trained for five iterations in each epoch. To extract features from the training and test images, all images were down-sampled to 256$\times$256 pixels to fit into the utilized GPU memory. For each experiment, an identical training set (70%), validation set (10%), and test set (20%) were used to ensure fair comparisons between different experiments. The data splitting was done randomly by keeping the ratio between different classes in the training, validation and test sets. In the process of graph construction, the parameter $k$ was set to 25 for the BreakHis dataset and 15 for the BACH dataset. For the attention layer in the A-ARVGAE model, a 2-head attention layer was used. All experiments were conducted on a single machine using an NVIDIA GTX 1060 GPU and an Intel Core™ i7-9750H CPU.
4.2 Experiments
We performed two sets of experiments for each dataset in this study. In the first set, we investigated the effect of using the CCL feature extractor against other pre-trained models for the breast cancer histology image retrieval task. The results of this experiment for the BreakHis and BACH datasets are reported in Table 2 and Table 3, respectively. For a fair comparison, we used the A-ARVGAE model in all experiments.
Table 2: Evaluation of the retrieval performance for the BreakHis dataset with different feature extractors. In all experiments attention-based adversarially regularized variational graph autoencoder is used as the retrieval model.
Magnification Feature Extraction mAP(5) mMV(5)
40×40$\times$40 × VGG19 0.945 0.859
MobileNetV2 0.957 0.894
DenseNet121 0.941 0.864
Xception 0.899 0.743
InceptionResNetV2 0.893 0.778
NASNetLarge 0.925 0.809
EfficientNetV2M 0.919 0.768
CCL 0.978 0.949
100×100$\times$100 × VGG19 0.948 0.850
MobileNetV2 0.947 0.884
DenseNet121 0.953 0.879
Xception 0.903 0.739
InceptionResNetV2 0.889 0.739
NASNetLarge 0.919 0.777
EfficientNetV2M 0.915 0.826
CCL 0.980 0.961
200×200$\times$200 × VGG19 0.941 0.840
MobileNetV2 0.955 0.870
DenseNet121 0.946 0.860
Xception 0.904 0.796
InceptionResNetV2 0.883 0.695
NASNetLarge 0.921 0.776
EfficientNetV2M 0.905 0.791
CCL 0.964 0.930
400×400$\times$400 × VGG19 0.945 0.861
MobileNetV2 0.941 0.850
DenseNet121 0.955 0.895
Xception 0.911 0.784
InceptionResNetV2 0.897 0.718
NASNetLarge 0.906 0.751
EfficientNetV2M 0.905 0.790
CCL 0.940 0.839
BreakHis Avg. VGG19 0.944 0.852
MobileNetV2 0.950 0.874
Xception 0.904 0.765
InceptionResNetV2 0.890 0.727
NASNetLarge 0.917 0.778
EfficientNetV2M 0.911 0.791
DenseNet121 0.948 0.874
CCL 0.965 0.919
Table 3: Evaluation of the retrieval performance for the BACH dataset with different feature extractors. In all experiments attention-based adversarially regularized variational graph autoencoder is used as the retrieval model.
Magnification Feature Extraction mAP(5) mMV(5)
200×200$\times$200 × VGG19 0.899 0.787
MobileNetV2 0.893 0.737
DenseNet121 0.897 0.700
Xception 0.887 0.687
InceptionResNetV2 0.872 0.675
NASNetLarge 0.892 0.800
EfficientNetV2M 0.882 0.762
CCL 0.947 0.913
As observed from Table 2 for the BreakHis dataset, the CCL feature extractor outperforms all other pre-trained feature extractor models across all metrics, except for 400$\times$ magnification, where DenseNet121 delivered better retrieval performance. However, for the average performance across all magnifications, CCL delivers the best results compared with all other pre-trained models, including DenseNet121. The results in Table 3 also indicate the superior performance of the CCL feature extractor compared with other pre-trained models for the BACH dataset.
Overall, the results reported in Table 2 and Table 3 for the BreakHis and BACH datasets show that the feature extractor trained on histological images outperforms feature extractors trained with non-medical images.
In the second experiment, we compare the performance of the employed A-ARVGAE model with other GNN-based and non-GNN-based models for both BreakHis and BACH datasets. The results of these experiments are presented in Table 4.
Table 4: Comparison of different models (GNN-based models including GAE, VGAE, and ARVGA, non-GNN-based model, and ours) for the BreakHis and BACH datasets for the breast cancer histological image retrieval task. GAE: graph autoencoder, VGAE: variational graph autoencoder, ARVGA: adversarial regularized variational graph autoencoder, RetCCL: retrieval with cluster-guided contrastive learning
Dataset Method mAP(5) mMV(5)
BreakHis (40$\times$) GAE 0.901 0.773
VGAE 0.962 0.934
ARVGA 0.969 0.909
RetCCL 0.983 0.9789
Ours 0.978 0.9499
BreakHis (100$\times$) GAE 0.910 0.792
VGAE 0.966 0.903
ARVGA 0.949 0.909
RetCCL 0.970 0.951
Ours 0.980 0.961
BreakHis (200$\times$) GAE 0.926 0.825
VGAE 0.955 0.900
ARVGA 0.949 0.879
RetCCL 0.920 0.920
Ours 0.964 0.930
BreakHis (400$\times$) GAE 0.924 0.828
VGAE 0.923 0.806
ARVGA 0.932 0.834
RetCCL 0.897 0.889
Ours 0.940 0.839
BreakHis Avg. GAE 0.915 0.804
VGAE 0.951 0.885
ARVGA 0.949 0.882
RetCCL 0.942 0.934
Ours 0.965 0.919
BACH (200$\times$) GAE 0.919 0.740
VGAE 0.937 0.798
ARVGA 0.935 0.817
RetCCL 0.901 0.8942
Ours 0.947 0.913
As shown in Table 4, the A-ARVGAE model surpasses all GNN-based models i.e., GAE, VGAE, and ARVGA for both datasets. The superior performance of the proposed model over ARVGA demonstrates the significance of the attention component in the employed architecture. An attention component allows the model to focus on specific parts of input data to calculate the model weights.
Compared with the RetCCL model, our approach delivers superior performance for both evaluation indices on the BACH dataset. However, for the BreakHis dataset, RetCLL shows slightly better performance for some magnifications. Nevertheless, our approach still yields superior results in other magnifications. The average results across all magnifications indicate the superior performance of our approach compared with RetCCL, based on the mAP(5) evaluation index (2.3% improved performance). However, for the MVP(5) score, RetCCL exhibits a better average performance with a 1.5% margin.
Overall, while our approach has demonstrated superior or very competitive performance in all experiments, it should be noted that GNN training typically requires large datasets [59]. The utilized model in this study was trained on the BreakHis and BACH datasets, which are relatively small.
For visual analysis, we randomly select five query images from the BreakHis and BACH test sets (four from the BreakHis with different magnifications and one from the BACH dataset) to visualize the model output (top five ranks). Figure 5 depicts the query images and model output. As the results show, the model is capable of providing relevant images (with identical classes as the query images) for most of the samples.
Query Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
BH (40$\times$)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
M M M M M M
BH (100$\times$)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
M BN M M M M
BH (200$\times$)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
M M M M BN BN
BH (400$\times$)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
M M M M M M
BACH (200$\times$)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
N N N N N N
Figure 5: Five sample queries from the BreakHis and BACH datasets and their top similar retrieved images. Successful and unsuccessful retrieval are denoted by () and (), respectively. M: Malignant class, BN: Benign class, N: Normal class, BH: BreakHis
As suggested in [60], we also compare the embedding vectors of some example image pairs with different degrees of similarity in Figure 6. As shown in the figure, the embedding vectors for similar image pairs are very close to each other. In contrast, the embedding vectors for dissimilar pairs are separable.
Figure 6: Comparison of embedding values for four image pairs. Similar pairs refer to images within the same class, and dissimilar pairs refer to images in different classes. The distances between the embedding of the two images in each image pair are shown on the right side of the figure. For each image pair, the left and right image embeddings are depicted in black and red, respectively.
5 Conclusion & Future Works
Recently, there has been a growing trend in utilizing GNNs to address challenges in analyzing histopathology images. This research introduces a new GNN-based approach specifically designed for breast histopathology image retrieval. We incorporated both attention mechanisms and adversarial training in a variational graph auto-encoder model to enhance retrieval performance. Moreover, instead of using conventional pre-trained models as the node feature extractor, we employed the CCL feature extractor, specifically pre-trained on histological images. Our results confirm that incorporating both attention mechanisms and adversarial training consistently exhibits superior performance compared to other GNN variants. Additionally, using the CCL feature extractor further boosts performance in our experiments. While our approach delivers promising results for breast histology image retrieval, certain aspects can be considered for further investigation in future studies.
Transformers-based models have shown impressive performance in various computer vision tasks. Some Transformer-based GNNs, such as graph transformer networks [61], have been developed for tasks such as node classification. Such models can be adapted and employed for the medical image retrieval task as well.
While we use FLANN to build the graph in our model, other state-of-the-art ANN methods and libraries, such as Faiss library  [62] or HNSW method  [8] can be exploited to enhance performance. However, it should be noted that most of these methods are slower in comparison to FLANN  [36].
Another potential enhancement for the model is replacing the inner product decoder with a non-linear neural network, a topic that can be addressed in future studies.
Finally, using well-known boosting techniques, such as test-time augmentation or ensembling, has been shown to enhance performance for tasks such as medical image classification or segmentation [63, 64, 65]. Such techniques can also be incorporated into the workflow, and their impact on the retrieval performance can be investigated.
Acknowledgments
This project is partially supported by the Ernst Mach Grant (reference number: MPC-2023-00569).
Conflict of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Credit Authorship Contribution Statement
Nematollah Saeidi: Conceptualization, Methodology, Investigation, Software, Writing – original draft. Hossein Karshenas: Supervision. Bijan Shoushtarian: Supervision. Sepideh Hatamikia: Supervision. Ramona Woitek: Supervision. Amirreza Mahbod: Conceptualization, Validation, Review & editing, Supervision.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used Grammarly (version 1.2.75) and ChatGPT (version 4) in order to check grammar and spelling and, at times, improve the readability of some sentences. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v3

@@
-Breast Histopathology Image Retrieval by Attention-based Adversarially Regularized Variational Graph Autoencoder with Contrastive Learning-Based Feature Extraction
+Leveraging Medical Foundation Model Features in Graph Neural Network-Based Retrieval of Breast Histopathology Images
 Abstract
@@
 Early detection and appropriate treatment can significantly reduce its impact.
-While histopathology examinations play a vital role in rapid and accurate diagnosis, they often require a substantial workforce and experienced medical experts for proper recognition and cancer grading.
+While histopathology examinations play a vital role in rapid and accurate diagnosis, they often require experienced medical experts for proper recognition and cancer grading.
 Automated image retrieval systems have the potential to assist pathologists in identifying cancerous tissues, thereby accelerating the diagnostic process.
-Nevertheless, due to considerable variability among the tissue and cell patterns in histological images, proposing an accurate image retrieval model is very challenging.
-This work introduces a novel attention-based adversarially regularized variational graph autoencoder model for breast histological image retrieval.
-Additionally, we incorporated cluster-guided contrastive learning as the graph feature extractor to boost the retrieval performance.
-We evaluated the performance of the proposed model on two publicly available datasets of breast cancer histological images and achieved superior or very competitive retrieval performance, with average mAP scores of 96.5% for the BreakHis dataset and 94.7% for the BACH dataset, and mVP scores of 91.9% and 91.3%, respectively.
+Nevertheless, proposing an accurate image retrieval model is challenging due to considerable variability among the tissue and cell patterns in histological images.
+In this work, we leverage the features from foundation models in a novel attention-based adversarially regularized variational graph autoencoder model for breast histological image retrieval.
+Our results confirm the superior performance of models trained with foundation model features compared to those using pre-trained convolutional neural networks (up to 7.7% and 15.5% for mAP and mMV, respectively), with the pre-trained general-purpose self-supervised model for computational pathology (UNI) delivering the best overall performance.
+By evaluating two publicly available histology image datasets of breast cancer, our top-performing model, trained with UNI features, achieved average mAP/mMV scores of 96.7%/91.5% and 97.6%/94.2% for the BreakHis and BACH datasets, respectively.
 Our proposed retrieval model has the potential to be used in clinical settings to enhance diagnostic performance and ultimately benefit patients.
 keywords:
-Computational pathology , Graph Neural Network , Graph Construction , Adversarial Training , Variational Graph Autoencoder , Attention Module , Image Retrieval
+Computational Pathology , Graph Neural Network , Foundation Models , Image Retrieval
 1 Introduction
@@
 According to 2018 statistics, breast cancer accounted for 15% of all cancer-related deaths and has increased by 24% in recent years [3].
-While breast cancer causes many deaths worldwide, early detection and diagnosis followed by appropriate treatment can help to reduce its impact [4].
+While breast cancer causes many deaths worldwide, early detection and diagnosis, followed by appropriate treatment, can help to reduce its impact [4].
 The need for early detection of breast cancer has led to new solutions, such as computer-aided diagnosis systems based on histological images [5].
@@
 However, proper diagnosis based on histological images often requires a substantial workforce and experienced medical experts.
-To overcome this challenge, pathologists can benefit from referring to archived databases for complex cases to find similar examples with verified findings to minimize diagnostic errors.
-Content-based image retrieval (CBIR) is an approach to identify visually similar images in a database and can be used for various applications [7, 8].
-For example, pathologists reviewing tissue sections for breast cancer diagnosis may benefit from searching through similar labeled histopathological image data to improve their diagnostic accuracy in distinguishing cancerous from non-cancerous tissue.
-CBIR techniques have also been used for various non-medical applications such as ”image search” for general and natural images, visual search for retail products, and multiple tools for recognizing faces and artworks [9].
-Prior machine learning-based CBIR systems often employed models adapted to specific applications, requiring the collection of labeled data for each particular use, which presented a significant implementation challenge.
-Additionally, the term ”similarity” in these approaches could have different interpretations depending on the specific applications [10].
-For instance, two histological images might be considered similar because they originate from the same organ, are classified as the same type of cancer, share similar staining, or represent comparable histological features.
-Content-based medical image retrieval (CBMIR), a subset of CBIR, is specifically adapted to medical images and has been used in the context of radiology and histology in former studies [9, 10, 11].
-CBMIR for histological images, not only aids histopathology in grading new tissue samples but also assists in analyzing patterns within similar tissues from previous patients [11, 5].
-CBMIR offers distinct advantages over classification tasks.
-For instance, in classifying prostate cancer, a classification model provides class labels for each input.
-In contrast, a CBMIR system produces the top-$k$ most similar images based on a query image, where $k$ represents the number of desired results.
-Upon availability, it can also provide other relevant information such as clinical meta-data or diagnosis reports [12].
-Therefore, CBMIR provides pathologists and practitioners with more comprehensive information about diverse tissue patterns and enables them to identify the most similar patterns within different tissues, which could potentially minimize intra-and inter-observer variability [9, 13].
-Another advantage of CBMIR is that it inherently includes a clinician-in-the-loop, which is often required for artificial intelligence (AI) applications in healthcare.
-Furthermore, a pathologist using CBMIR in the diagnostic process is required to come to their own diagnostic conclusion after incorporating additional information provided by CBMIR systems rather than replacing the pathologist’s diagnosis with an AI-based classifier.
-Unlike histological images with their reports written by pathologists that are presented by CBMIR systems, a classifier is not necessarily easily interpretable and - even if presented to a clinician as a second opinion - may lead to automation bias [14, 15].
-From a machine learning perspective, CBMIR is typically performed using unsupervised techniques, while classifiers are usually trained with supervised methods that require labeled datasets.
-Histopathology image retrieval is a very challenging task due to the substantial tissue and cell pattern variability [10].
+Pathologists can benefit from referring to archived databases for complex cases to find similar examples with verified findings to minimize diagnostic errors and overcome this challenge.
+Content-based image retrieval (CBIR) is an approach to identifying visually similar images in a database and can be used for various applications [7, 8].
+Content-based medical image retrieval (CBMIR), a subset of CBIR, is specifically adapted to medical images and has been used in the context of radiology and histology in former studies [9, 10].
+The CBMIR of histological images not only aids histopathology in grading new tissue samples but also assists in analyzing patterns within similar tissues from previous patients [10, 5].
+Histopathology image retrieval is challenging due to the substantial tissue and cell pattern variability [11].
 Various semi- and fully-automatic approaches have been proposed for histopathology image retrieval [9].
-Among these, graph neural networks (GNN) have gained considerable attention in recent years.
-Considering the grid data structure in images and the tremendous developments in convolutional neural networks (CNNs), many graph convolutional neural network (GCNN)-based architectures have also been extensively proposed for a wide range of computer vision tasks, including image retrieval [16].
-In GCNN, input images can be represented as the graph nodes, and their features are usually extracted by pre-trained CNNs such as Visual Geometry Group (VGG) [17] or EfficientNet [18] to construct the graph.
-However, the exploited pre-trained models are typically trained on natural images, and their extracted features may not be optimally suited for medical data (e.g., histological images) [19, 20].
+Among these, graph neural networks (GNN) and graph convolutional neural networks (GCNN) have gained considerable attention in recent years [12].
 While supervised learning is usually used to train GCNNs, some GCNN variants, such as variational graph autoencoder (VGAE) [13], can be used for unsupervised learning tasks.
-In VGAE, the encoder network maps input data to a lower-dimensional latent space, and the decoder network maps latent space vectors back to the original input data.
-Unlike conventional graph autoencoders (GAE), VGAE utilizes probabilistic encoding that maps input data to a probability distribution over the latent space.
 VGAE performance can be further improved by incorporating other techniques, such as adversarial training or adding an attention mechanism to the model [13].
-In adversarially regularized variational graph autoencoder (ARVGA) [22], the encoder is forced to generate embedding vectors resembling the data from a prior distribution through adversarial training.
-Incorporating an attention mechanism into CNNs [23] or GNNs [24] is another technique that has been exploited.
+In an adversarial regularized variational graph autoencoder (ARVGA) [14], the encoder is forced to generate embedding vectors resembling the data from a prior distribution through adversarial training.
+Incorporating an attention mechanism into convolutional neural networks (CNNs) [15] or GNNs [16] is another technique that has been exploited.
 These techniques have also been jointly applied for specific applications such as graph link prediction [17], graph clustering visualization [17], and node classification in partially labeled graphs [18].
-In this study, for the first time, we adapted and exploited attention-based adversarially regularized variational graph autoencoders (A-ARVGAE) for the histological breast image retrieval task.
-Moreover, unlike previous studies that use conventional pre-trained CNNs for feature extraction to build the graph, we used the recently developed clustering-guided contrastive learning (CCL) feature extractor [27], which is explicitly designed for histopathological images.
-We evaluated the performance of our proposed approach on two binary and multi-class histopathological breast image datasets, namely breast cancer histopathological database (BreakHis) [28] and breast cancer histology images dataset (BACH [29]), and achieved superior or very competitive performance compared to other state-of-the-art models.
+Generally in recent studies based on GNNs for image retrieval, input images are represented as the graph nodes whose features are usually extracted by pre-trained CNNs to construct the graph.
+However, the exploited pre-trained CNN models are typically trained on natural images based on the ImageNet dataset [19], and their extracted features may not be optimally suited for medical data (e.g., histological images) [20, 21].
+On the other hand, recently, several pre-trained foundation models, trained on millions of data samples in an unsupervised manner, have been developed and applied to a wide range of computer vision tasks.
+In this study, we adapted and exploited attention-based adversarially regularized variational graph autoencoders (A-ARVGAE) for the histological breast image retrieval task.
+Unlike most prior studies that employ conventional pre-trained CNNs for feature extraction to construct the graph, we used medical foundation models, including contrastive language-image pre-training with medical data (BioMedCLIP) [22], general-purpose self-supervised model for computational pathology (UNI) [23], and clustering-guided contrastive learned model (CCL) [24], as the graph node feature extractors.
+Evaluated on two binary and multi-class histopathological breast image datasets, namely the breast cancer histopathological database (BreakHis) [25] and the breast cancer histology image dataset (BACH) [26], our results demonstrate the superior performance of models trained with foundation model features over those with CNN features, with UNI delivering the best overall retrieval performance.
 The main contribution of this research can be summarized as follows:
   • 1.
-    We adapted and exploited one of the state-of-the-art GNN models, which incorporated both attention mechanism and adversarial training for breast cancer histology image retrieval task.
+    We adapted and exploited one of the state-of-the-art GNN models, which incorporated both attention mechanism and adversarial training for the histology image retrieval task of breast cancer.
   • 2.
-    Instead of using conventional pre-trained CNNs as the node feature extractor, we made use of CLL, which is explicitly designed for histological images.
+    Instead of using conventional pre-trained CNN features as the node feature extractor, we made use of medical foundation model features.
   • 3.
-    Through extensive experiments, we show the excellent performance of our approach for both utilized GNN and feature extraction on two histopathological breast image datasets, namely BreakHis and BACH, and achieved superior or very competitive performance compared to other approaches.
-The rest of this paper is organized as follows: In Section 2, we present a background on methods for medical image retrieval and graph representation learning in histopathology.
+    Through extensive experiments, we show the excellent performance of our proposed approach for both utilized GNN and feature extraction on two histopathological breast image datasets.
+The rest of this paper is organized as follows: In Section 2, we present a brief background on methods for medical image retrieval and graph representation learning in histopathology.
 Section 3 describes our methodology, which contains all steps for the proposed framework.
 Section 4 consists of the experiments and evaluations with an analysis of the results.
-Finally, in Section 5, we conclude the paper with suggestions for future works.
+Finally, Section 5 concludes the paper with suggestions for future works.
 2 Related Works
-This section contains a literature review for medical image retrieval and graph representation learning in histopathology, considering their relevance to our work.
+This section contains a brief review of medical image retrieval and graph representation learning in histopathology, considering their relevance to our work.
 2.1 Medical Image Retrieval
-Various methods have been proposed in the literature for medical image retrieval.
-Text-based image retrieval systems were among the first attempts to find similar images in a given database.
-However, these methods rely heavily on expert annotation and proper keyword selection, which makes them generally considered ineffective and time-consuming [30].
-CBMIR models, on the other hand, rely on the content (features) of the images instead of their given labels.
-Consequently, they can be used in an unsupervised manner without the need for manual labeling.
+Various methods have been proposed in the literature for CBMIR.
+CBMIR models rely on the content (features) of the images and can be used unsupervised without the need for manual labeling.
 Conventional CBMIR models consist of image preprocessing, hand-crafted or automatic feature extraction, and feature similarity measurements using different distance-based methods.
-To enhance interpretability, some approaches propose reducing the disparity between the extracted features and the high-level semantic meanings of the images by incorporating expert feedback or medical metadata into the retrieval model [30, 31].
-Commonly used preprocessing methods include noise reduction, contrast enhancement, patch generation (e.g., from whole slide images), and normalization (e.g., stain normalization in histological images) [32, 33].
+Commonly used preprocessing methods include noise reduction, contrast enhancement, patch generation from whole slide images (WSIs), and normalization (e.g., stain normalization in histological images) [27, 28].
 However, many novel machine learning (ML)- and deep learning (DL)-based feature extractors do not require any preprocessing steps.
-For feature extraction, both manually designed and DL-based automatic feature extractors have been utilized in former studies.
-Hand-crafted features are employed to depict the morphological and textural characteristics of images.
-These features include color histogram-based features, histogram of oriented gradients (HOG) descriptor, local statistical features, global image structure tensor (GIST) features, Gabor features, wavelet-based features, intensity-based features, textural features, gray level co-occurrence matrix (GLCM) features, and scale-invariant feature transform (SIFT) descriptors [30, 27].
-It should be noted that experts usually need to set up specific transformation functions to derive these features.
-However, more recently, the rapidly evolving field of DL has made it easier to automatically extract high-level features from the images [10].
-Many pre-trained DL-based models, such as VGG19, ResNet, MobileNetV2, Xception, EfficientNet, and DenseNet121, have been used as feature extractors for various computer vision tasks such as medical image classification or medical image retrieval [34, 35].
-Hand-crafted or automatically extracted features of a query image are then compared to the image features in a database to find similar images.
+For feature extraction, both manually designed and more recently DL-based automatic feature extractors have been utilized.
+Many pre-trained DL-based models, such as EfficientNet [29], and DenseNet121 [30], have been employed as feature extractors for various computer vision tasks including medical image retrieval [31, 32].
+Hand-crafted or automatically extracted features of a query image are then compared to the image features in a database to find similar ones.
 For this comparison, various similarity measurements such as distance-based retrieval (SDR), customized query approach (CQA), or nearest-neighbor matching have been used [33].
-In the context of histopathology, Yottixel [37] and FISH [13] methods select the most distinctive patches to represent whole slide images (WSIs) through $K$-means clustering based on RGB histogram and spatial coordinate features.
-Yottixel employs a pre-trained DenseNet121, trained on the ImageNet dataset [38], to extract patch embeddings, which are then transformed into binary barcodes.
+In the context of histopathology, Yottixel [34] and FISH [35] methods select the most distinctive patches to represent WSIs through $K$-means clustering based on RGB histogram and spatial coordinate features.
+Yottixel employs a pre-trained DenseNet121, trained on the ImageNet dataset, to extract patch embeddings, which are then transformed into binary barcodes.
 FISH trains a model based on an autoencoder on the cancer genome atlas (TCGA) dataset to encode an integer index for each patch during retrieval.
-However, it is worth mentioning that using a feature extractor trained only on natural images might be sub-optimal due to the fundamental differences between natural and histopathological images.
-In contrast, CCL is a newly developed model specifically designed for histological images.
-CCL pretraining was performed with around 15 million image patches extracted from more than 32,000 WSIs.
-The CCL learning method utilizes large-scale unlabeled histopathological images for universal feature extraction that can be directly applied for various WSI analysis tasks such as retrieval.
-In our study, we utilized CCL as a robust and relevant feature extractor to build the graph.
+However, it is worth mentioning that using a feature extractor trained only on natural images based on the ImageNet dataset might be sub-optimal due to the fundamental differences between natural and medical/histopathological images.
+With the advent of transformer-based models, the availability of large-scale unlabeled data and advanced vision-based or vision-language-based training methods such as contrastive language-image pre-training (CLIP) [36] or knowledge distillation with no Labels (DINO) [37], several foundation models have been developed and pre-trained on large scale data in an unsupervised or self-supervised manner.
+These foundation models have been used for various applications such as off-the-shelf feature extractors for computer vision tasks and include general-purpose foundation models such as OpenCLIP or medical foundation models such as UNI (for histopathological images) [23] or PanDerm (for dermatoscopic images) [38].
+In this work, in addition to using well-known CNN model features for benchmarking, we employed three medical foundation models as graph node feature extractors and compared their results.
+The foundation models utilized include BioMedCLIP (a medical foundation model trained on diverse medical data types) and UNI and CCL (where both are medical foundation models specifically trained on histological images).
 2.2 Graph Representation Learning in Histopathology
@@
 In contrast to CNNs, GNNs are better at preserving complex correlations among neighboring elements during the learning process.
-GNNs have been successfully applied to various tasks, including tumor/object localization, classification, survival prediction, and image retrieval [39].
 In the field of computational pathology, several GNN models, such as cell graphs, tissue graphs, cell-tissue graphs, and patch graphs, have been proposed for tasks like node classification (e.g., nuclei classification) or graph classification (e.g., histologic image patch categorization) [39, 40].
@@
 Using other models like HoVerNet [41] or DDU-Net [42] to generate nuclei segmentation masks as input for the graph-based model is an example of such preprocessing.
-Regarding clinical applications, GNNs have been used in different cancer types analysis, including breast cancer, colorectal cancer, and prostate cancer, among others [41].
-To build a graph, one needs to define nodes, node embeddings, and edges.
-In the context of histological images, nodes could represent nuclei, image patches, or tissue regions.
-Node embeddings are features of the nodes that can be extracted either manually or automatically using pre-trained models or other techniques.
-The edges depict connections between nodes, typically determined by predefined criteria such as $k$-nearest neighbors or distances between nodes.
-Once the graph is built, various GNN or GCN models can be applied, depending on the specific applications.
-A GCN contains several key components, including an aggregation layer to define how features from neighbor nodes are aggregated, an activation layer to incorporate non-linearity into the model, graph pooling to reduce graph size (if needed), and graph convolutional operations to update node representations.
-It’s worth noting that other variants, such as GraphSAGE [44], GAT [45], or ChebNet [46] have been introduced in the literature as novel propagation rules to enhance performance across different applications.
-The output of a GNN typically depends on the application.
-For graph classification, a multilayer perceptron (MLP) can be added to the graph.
-In the case of graph reconstruction for unsupervised tasks, a linear projection decoder can be added to the model.
 As highlighted in a recent review [43], most GNN-based works in computational pathology are related to supervised classification tasks.
@@
 However, even in these studies, the potential of attention mechanisms and adversarial training have not been explored.
-In our research, for the first time, we adapted and made use of both attention mechanism and adversarial training in variational graph autoencoders for the histological image retrieval task.
+In our research, for the first time, we adapted and used both attention mechanism and adversarial training in variational graph autoencoders for the histological image retrieval task.
 3 Methodology
-The general workflow of our proposed approach for histological image retrieval is depicted in Figure 1. In the following, we discuss the different parts of the workflow in detail.
-Figure 1: The general workflow of the proposed breast histological image retrieval model involves distinct processes for training and testing.
+The general workflow of our proposed approach for histological image retrieval is depicted in Fig. 1. In the following, we discuss the different parts of the workflow in detail.
+Figure 1: The proposed general workflow of the breast histological image retrieval model involves distinct processes for training and testing.
 A-ARVGAE: Attention-based adversarially regularized variational graph autoencoders.
 3.1 Dataset
-The dataset refers to a collection of histological image patches used to train and test the model.
-Each image has certain features and characteristics related to a specific class or histopathological grade.
-We use two breast histopathology image datasets in this study, namely the BreakHis dataset [28] and the BACH dataset [29], which are subsequently detailed.
-Example images from the datasets are shown in Figure 2.
+We use two breast histopathology image datasets in this study, namely the BreakHis and the BACH dataset, which are subsequently detailed. Example images from the datasets are shown in Fig. 2.
   • 1.
     BreakHis: This dataset contains hematoxylin and eosin (H&E)-stained breast histology microscopy images captured at various magnifications (40×40$\times$40 ×, 100×100$\times$100 ×, 200×200$\times$200 ×, 400×400$\times$400 ×) and is categorized into two primary classes, including benign and malignant classes.
-The dataset contains 7,909 images including 2,480 benign and 5,429 malignant samples with a size of $700\times 460$ pixels in Portable Network Graphics (PNG) format.
+The dataset contains 7,909 images of which are 2,480 benign and 5,429 malignant samples with a size of $700\times 460$ pixels in Portable Network Graphics (PNG) format.
   • 2.
@@
 3.2 Feature Extraction
-The first step in our proposed retrieval model is the extraction of features for all images in the dataset.
-Instead of using hand-crafted features or conventional pre-trained CNNs as feature extractors, we utilize CCL as the main feature extractor.
-Unlike other conventional pre-trained CNNs, CCL is trained with histological images.
-We hypothesize that using such a pre-trained feature extractor could lead to better descriptive features for the task of histological image retrieval.
-We also use a set of standard baseline models as feature extractors and compare their retrieval performance with the CCL feature extractor.
-For this, we employ well-known pre-trained models, namely VGG19 [17], MobileNetV2 [48], DenseNet121 [49], Xception [50], InceptionResNetV2 [51], NASNetLarge [52], and EfficientNetV2M [18].
-The details of these baseline pre-trained models are presented in Table 1.
-It is worth mentioning that all these pre-trained models are initially trained on natural images using the ImageNet dataset [38], but the CCL feature extractor used in this study is pre-trained with histological image patches.
-The results of these comparisons are reported in Section 4 (Table 2 and Table 3).
-Table 1: Details of baseline feature extractors. All these pre-trained models are initially trained on natural images using the ImageNet dataset [38].
-Feature Extraction Model Size (MB) Parameters Depth
-VGG19 549 143.7 M 19
-MobileNetV2 14 3.5 M 105
-DenseNet121 33 8.1 M 242
-Xception 88 22.9 M 81
-InceptionResNetV2 215 55.9 M 449
-NASNetLarge 343 88.9 M 533
-EfficientNetV2M 220 54.4 M -
+The first step in our proposed retrieval model is the feature extraction of all images in the dataset.
+Instead of using hand-crafted features or conventional pre-trained CNNs as feature extractors, we utilize pre-trained medical foundation models as the main feature extractors.
+Unlike conventional pre-trained CNNs, the foundation models in this study are specifically trained on medical/histological images.
+We leverage three medical foundation models in this study, namely BioMedCLIP, UNI and CCL.
+BioMedCLIP is a vision-language model based on the CLIP architecture and is initially trained on the PMC15M dataset, which contains approximately 15 million image-text pairs extracted from PubMed medical articles.
+It utilizes a vision transformer (ViT) as its image encoder and a bidirectional encoder representation from transformers (BERT) model as its text encoder.
+In this study, we use only the image encoder for feature extraction.
+UNI is a vision model trained with a self-supervised DINO V2 [46] approach on over 100 million histological image patches sourced from 100,000 WSIs.
+UNI uses ViT-based models in student and teacher branches and has shown superb performance across various computer vision tasks.
+CCL is also a vision model, trained using a novel self-supervised, clustering-guided contrastive learning technique on over 15 million histological image patches.
+The CCL approach integrates two contrastive losses (a weighted InfoNCE and a group-level InfoNCE).
+It utilizes a clustering-guided memory bank to optimize positive and negative sample identification during contrastive learning.
+Further details on these models can be found in their respective studies [22, 23, 24].
+We hypothesize that these models, trained on vast datasets of medical images, will yield superior features for histological image retrieval.
+We also use standard baseline models as feature extractors and compare their retrieval performance with the foundation model feature extractors.
+For this purpose, we employ well-known pre-trained CNN models, namely VGG19 [47], MobileNetV2 [48], DenseNet121 [30], NASNetLarge [49], and EfficientNetV2M [29].
 3.3 Graph Construction
-GNNs are mostly applied on graph-based datasets.
+GNNs are mainly applied to graph-based datasets.
 These datasets typically include nodes, node features, edges, and adjacency matrices.
@@
 This makes graph construction an essential step for applying GNNs to such data.
-For the retrieval task, graph construction methods often treat subjects (e.g., images) as graph nodes ($V$), with the features of these subjects acting as node features ($X$).
-Edges between nodes ($E$) can be defined using distance metrics like cosine and Euclidean distances, indicating the likelihood of interaction between entities [53, 54, 55, 56].
-The topology of these graphs is usually determined experimentally, employing strategies such as a pre-defined proximity threshold, a nearest neighbor rule, or a probabilistic model.
-To create edges between nodes, various algorithms such as $k$ approximate nearest neighbor ($k$-ANN) can be used.
-This approach is more time-efficient compared to the $k$-NN algorithm and is beneficial, especially in high-dimensional feature spaces, where the complexity of high dimensions slows down exact nearest neighbor searches.
-While there are various implementations of ANN algorithms, the ANN benchmark [57] is a widely-used library that encompasses most ANN-based methods.
-In our study, we utilize the fast library for approximate nearest neighbors (FLANN) to calculate distances between node embeddings, selecting the $k$ nearest nodes to each entity as adjacent nodes.
-The computation of graph edges is conducted as follows:
-$e_{ij}=\begin{cases}1&\text{if }v_{j}\in k\text{-ANN}(v_{i})\\ 0&\text{otherwise}\end{cases}\quad(1)$
-where $e_{ij}=1$ indicates the existence of an edge connection between the $i$-th and $j$-th nodes in the graph, and $e_{ij}=0$ denotes no edge.
-Utilizing the edges and nodes, the adjacency matrix ($A$) can be constructed where the values of the matrix ($a_{ij}$) act as indicators, quantifying of the connections between nodes $V_{i}$ and $V_{j}$.
-It is worth mentioning that we use transductive learning, where the entire dataset (training, validation, and test data) is observed to build the graph.
-The summarized workflow to build the graph in our study is shown in Figure 3.
-Figure 3: The framework of graph construction. CCL: Cluster guided contrastive learning, $K$-ANN: $K$ approximate nearest neighbors, FLANN: fast library for approximate nearest neighbors
+For the retrieval task, graph construction methods often treat subjects (e.g., images) as graph nodes, with the features of these subjects acting as node features.
+Edges between nodes can be defined using distance metrics like cosine and Euclidean distances, indicating the likelihood of interaction between entities [50, 51].
+By utilizing the edges and nodes, the adjacency matrix can be constructed where the values of the matrix act as indicators, quantifying the connections between nodes.
+The topology of the graphs is usually determined experimentally, employing strategies such as a pre-defined proximity threshold, a nearest neighbor rule, or a probabilistic model.
+Various algorithms, such as $k$ approximate nearest neighbor ($k$-ANN), can create edges between nodes.
+This approach is more time-efficient than the $k$-NN algorithm and beneficial, especially in high-dimensional feature spaces, where the complexity of high dimensions slows down exact nearest neighbor searches.
+While there are various implementations of ANN algorithms, the ANN benchmark [52] is a widely used library encompassing most ANN-based methods.
+In this study, we utilize the fast library for approximate nearest neighbors (FLANN) to calculate distances between node embeddings, selecting the $k$-nearest nodes to each entity as adjacent nodes.
+We use transductive learning, where the entire dataset (training, validation, and test data) is observed to build the graph. The summarized workflow of building the graph is shown in Fig. 3.
+Figure 3: The framework of graph construction. $K$-ANN: $K$ approximate nearest neighbors, FLANN: fast library for approximate nearest neighbors
 3.4 A-ARVGAE
-The overall architecture of the A-ARVGAE model is depicted in Figure 4.
-The proposed architecture comprises four main components: the encoder (yellow block), which also serves as the generator for adversarial training; the embedding latent space (pink block); the decoder (blue block); and the discriminator (green block), which is another essential component for adversarial training.
+The overall architecture of the A-ARVGAE model is depicted in Fig.
+4.
+The proposed architecture comprises four main components: the encoder (yellow block), which also serves as the generator for adversarial training, the embedding latent space (pink block), the decoder (blue block), and the discriminator (green block), which is another essential component for adversarial training.
 The constructed graph, as described in the previous section, is fed into the encoder of the A-ARVGAE model.
 The encoder includes a graph attention layer that performs attention operations, batch normalization, leaky ReLU, and dropout, followed by two graph convolutional layers.
-The attention-based convolutional operation integrating attention mechanisms into convolutional layers assigns varying weights to the neighbors, enabling the model to concentrate on the most relevant parts of the graph when learning node representations [45].
-The attention mechanism aggregates information from adjacent nodes with an adaptive weight allocation to neighboring nodes.
-As indicated by prior studies, the use of a multi-head attention layer has been shown to yield superior performance [25] and hence we also employ two-head attention layer in our work.
+The attention-based convolutional operation integrates attention mechanisms into convolutional layers by assigning varying weights to the neighbors and enabling the model to concentrate on the most relevant parts of the graph when learning node representations [53].
+As indicated by prior studies, a multi-head attention layer has yielded superior performance [17]; hence, we also employ the two-head attention layer in our work.
 Inspired by the VGAE [20], the encoder in our architecture does not encode an input into a fixed latent representation.
-Instead, the outputs of the graph convolutional layers are parameters (mean ($\mu$) and variance ($\sigma$)) that define a probability distribution over the possible latent representations, denoted as $Z$, characterized by a normal distribution $N(\mu,\sigma)$.
+Instead, the outputs of the graph convolutional layers are parameters (mean ($\mu$) and variance ($\sigma$)) that define a probability distribution over the possible latent representations, denoted as $Z$, characterized by the Normal distribution $N(\mu,\sigma)$.
 Sampling from the latent space involves drawing samples from the distribution $N(\mu,\sigma)$ to produce latent node embeddings.
@@
 The discriminator block comprises a straightforward MLP with two linear layers along with batch normalization, leaky reLU, and dropout to distinguish between real (a prior Gaussian distribution) and fake (latent space $Z$) embeddings.
-The complete model integrates three loss components: a regression-based reconstruction loss (related to VGAE), a Kullback–Leibler divergence loss (related to VGAE), and a binary cross-entropy loss (related to discriminator in adversarial training).
+The complete model integrates three loss components: a regression-based reconstruction loss (associated with VGAE), a Kullback–Leibler divergence loss (associated with VGAE), and a binary cross-entropy loss (related to discriminator in adversarial training).
 Further details regarding the model architecture and cost functions can be found in the respective studies [14, 53].
-We compare the results of our employed GNN model (A-ARVGAE) with other GNN-based models, namely GAE, VGAE, and ARVGA, and also with another non-GNN-based state-of-the-art approach, called RetCCL, represented in [27].
-RetCCL is a model that is recently developed for the histological image retrieval task, which is based on CCL feature extraction and nearest neighbor search to find similar images in a given dataset.
+As an ablation study, we compare the results of our employed GNN model (A-ARVGAE) with other GNN-based models, namely GAE, VGAE, and ARVGA.
 The results of these comparisons are reported in Section 4 (Table 4).
-Figure 4: Architecture of attention-based adversarially regularized variational graph autoencoder (A-ARVGAE).
+Figure 4: The Architecture of attention-based adversarially regularized variational graph autoencoder (A-ARVGAE).
 For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
@@
 The embeddings generated by the trained A-ARVGAE encoder in the previous step for each training image are stored in a database.
-This database is then utilized in the testing phase to identify the top-$K$ similar images for a given query image as explained in Section 3.7
+This database is then utilized in the testing phase to identify the top-$K$ similar images for a given query image as explained in Section 3.7.
 3.7 Testing phase
-As depicted in the lower part of Figure 1, for a query image, the model aims to identify the top $K$ similar images in the database.
-Similar to the process with training images, features are first extracted from the test image using the CCL feature extractor.
+As depicted in the lower part of Fig.
+1, the model aims to identify similar top-$K$ images in the database for a query image.
+Similar to the process with training images, features are first extracted from the test image using the pre-trained feature extractors.
 These features are then fed into the trained A-ARVGAE model to obtain the embedding in a lower-dimensional space, specifically from the encoder part of the model.
@@
 To find the most similar images, the ranking algorithm evaluates and ranks the candidate images based on Euclidean distance.
-These selected images are then arranged and ranked in ascending order according to Euclidean distances.
-Once ranked, the top-$K$ similar images are chosen as the final output of the retrieval model.
-To evaluate the performance of the developed approach, we use two evaluation metrics, namely, mean Average Precision (mAP) [58] and mean Majority Vote (mMV) [13].
-mAP is the average of the precision values computed at the ranks of relevant images in the retrieved results.
-It is a widely used metric for evaluating the effectiveness of a retrieval system.
-The formula for mAP is as follows:
-$mAP(k)=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\frac{1}{|R_{q}|}\sum_{k=1}^{|R_{q}|}\text{Precision}(k)\cdot\text{rel}(k)\quad(17)$
-Where $|Q|$ is the number of queries, $R_{q}$ is the number of relevant images for query image $q$, $\text{Precision}(k)$ is the precision at rank $k$ (the ratio of relevant images retrieved in the top $k$ results to the total number of retrieved images up to rank $k$), $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise.
-||$\cdot$|| ⋅ | denotes the length of the set.
-mMV evaluates the frequency with which the primary diagnosis among the top-$k$ results aligns with the queried diagnosis as follows:
-$mMV(k)=\frac{1}{|Q|}\sum_{q=1}^{|Q|}1\{L_{q}\in MV(\text{rel}(k))\}\quad(18)$
-Where $L_{q}$ represents the ground truth diagnosis of the image, and MV is the predicted diagnosis of an image taken from the majority vote of top-$k$ samples.
-Similar to mAP, $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise.
-Here, ||$\cdot$|| ⋅ | also denotes the length of the set.
+These images are then arranged and ranked in ascending order according to Euclidean distances.
+Once ranked, similar top-$K$ images are chosen as the final output of the retrieval model.
+To evaluate the performance of the developed approach, we use two widely-employed evaluation metrics for image retrieval, namely, mean Average Precision (mAP ($k$)) [54] and mean Majority Vote (mMV ($k$)) [35], where $k$ represents the top k results.
 4 Results & Discussion
@@
 4.1 Implementation Details
-We used the PyTorch deep learning framework (version 2.2.0) and the PyTorch Geometric library (version 2.4.0) to implement the models.
+We implemented the graph neural network using the PyTorch deep learning framework and the PyTorch Geometric library.
+We used the Keras framework for feature extraction with pre-trained CNN models, where the feature extraction from BioMedCLIP, UNI, and CCL models was performed using their respective repositories on Hugging Face and GitHub  [^1] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224
+https://github.com/mahmoodlab/UNI
+https://github.com/Xiyue-Wang/RetCCL
 Adam optimizer was used to train the models with an initial learning rate (LR) of 0.0001.
 For dropout layers, a probability of 0.2 was chosen in the model architecture.
-All models were trained for 250 epochs, whereas in models with adversarial components, the discriminator was trained for five iterations in each epoch.
-To extract features from the training and test images, all images were down-sampled to 256$\times$256 pixels to fit into the utilized GPU memory.
+All models were trained for 250 epochs, but for the models with adversarial components, the discriminator was trained for five iterations in each epoch.
+All images were down-sampled to 256$\times$256 pixels to fit into the utilized GPU memory to extract features from the training and test images.
 For each experiment, an identical training set (70%), validation set (10%), and test set (20%) were used to ensure fair comparisons between different experiments.
-The data splitting was done randomly by keeping the ratio between different classes in the training, validation and test sets.
-In the process of graph construction, the parameter $k$ was set to 25 for the BreakHis dataset and 15 for the BACH dataset.
+The data splitting was done randomly by keeping the ratio between different classes in the training, validation, and test sets.
+In the graph construction process, the parameter $k$ was set to 25 for the BreakHis dataset and 15 for the BACH dataset.
 For the attention layer in the A-ARVGAE model, a 2-head attention layer was used.
@@
 We performed two sets of experiments for each dataset in this study.
-In the first set, we investigated the effect of using the CCL feature extractor against other pre-trained models for the breast cancer histology image retrieval task.
-The results of this experiment for the BreakHis and BACH datasets are reported in Table 2 and Table 3, respectively.
+In the first set, we investigated the effect of using the medical foundation model feature extractors against other pre-trained CNN models for the histology image retrieval task for breast cancer.
+The results of this experiment for the BreakHis and BACH datasets are reported in Table 1, Table 2, and Table 3.
+For the BreakHis dataset, Table 1 shows the results per image magnification, while Table 2 provides the average results across magnifications.
 For a fair comparison, we used the A-ARVGAE model in all experiments.
-Table 2: Evaluation of the retrieval performance for the BreakHis dataset with different feature extractors.
-In all experiments attention-based adversarially regularized variational graph autoencoder is used as the retrieval model.
+Table 1: Evaluation of the retrieval performance for the BreakHis dataset with different feature extractors across different magnifications.
+The best results for convolutional neural networks and foundation models are highlighted by underlining and bold font, respectively.
+The attention-based adversarially regularized variational graph autoencoder is used in all experiments as the retrieval model.
 Magnification Feature Extraction mAP(5) mMV(5)
@@
 DenseNet121 0.941 0.864
-Xception 0.899 0.743
-InceptionResNetV2 0.893 0.778
 NASNetLarge 0.925 0.809
 EfficientNetV2M 0.919 0.768
+BioMedCLIP 0.983 0.934
+UNI 0.985 0.954
 CCL 0.978 0.949
@@
 DenseNet121 0.953 0.879
-Xception 0.903 0.739
-InceptionResNetV2 0.889 0.739
 NASNetLarge 0.919 0.777
 EfficientNetV2M 0.915 0.826
+BioMedCLIP 0.979 0.951
+UNI 0.973 0.946
 CCL 0.980 0.961
@@
 DenseNet121 0.946 0.860
-Xception 0.904 0.796
-InceptionResNetV2 0.883 0.695
 NASNetLarge 0.921 0.776
 EfficientNetV2M 0.905 0.791
+BioMedCLIP 0.968 0.925
+UNI 0.975 0.910
 CCL 0.964 0.930
@@
 DenseNet121 0.955 0.895
-Xception 0.911 0.784
-InceptionResNetV2 0.897 0.718
 NASNetLarge 0.906 0.751
 EfficientNetV2M 0.905 0.790
+BioMedCLIP 0.935 0.861
+UNI 0.938 0.850
 CCL 0.940 0.839
+Table 2: Evaluation of the average retrieval performance for the BreakHis dataset with different feature extractors.
+The best results for convolutional neural networks and foundation models are highlighted by underlining and bold font, respectively.
+The attention-based adversarially regularized variational graph autoencoder is used in all experiments as the retrieval model.
+Magnification Feature Extraction mAP(5) mMV(5)
 BreakHis Avg. VGG19 0.944 0.852
 MobileNetV2 0.950 0.874
-Xception 0.904 0.765
-InceptionResNetV2 0.890 0.727
+DenseNet121 0.948 0.874
 NASNetLarge 0.917 0.778
 EfficientNetV2M 0.911 0.791
-DenseNet121 0.948 0.874
+BioMedCLIP 0.966 0.917
+UNI 0.967 0.915
 CCL 0.965 0.919
 Table 3: Evaluation of the retrieval performance for the BACH dataset with different feature extractors.
-In all experiments attention-based adversarially regularized variational graph autoencoder is used as the retrieval model.
+The best results for convolutional neural networks and foundation models are highlighted by underlining and bold font, respectively.
+The attention-based adversarially regularized variational graph autoencoder is used in all experiments as the retrieval model.
 Magnification Feature Extraction mAP(5) mMV(5)
@@
 DenseNet121 0.897 0.700
-Xception 0.887 0.687
-InceptionResNetV2 0.872 0.675
 NASNetLarge 0.892 0.800
 EfficientNetV2M 0.882 0.762
+BioMedCLIP 0.942 0.846
+UNI 0.976 0.942
 CCL 0.947 0.913
-As observed from Table 2 for the BreakHis dataset, the CCL feature extractor outperforms all other pre-trained feature extractor models across all metrics, except for 400$\times$ magnification, where DenseNet121 delivered better retrieval performance.
-However, for the average performance across all magnifications, CCL delivers the best results compared with all other pre-trained models, including DenseNet121.
-The results in Table 3 also indicate the superior performance of the CCL feature extractor compared with other pre-trained models for the BACH dataset.
-Overall, the results reported in Table 2 and Table 3 for the BreakHis and BACH datasets show that the feature extractor trained on histological images outperforms feature extractors trained with non-medical images.
-In the second experiment, we compare the performance of the employed A-ARVGAE model with other GNN-based and non-GNN-based models for both BreakHis and BACH datasets.
+As observed from Table 1, for the BreakHis dataset, models trained with foundation model features outperform all models trained with CNN features across both metrics, except for 400$\times$ magnification, where DenseNet121 delivered better retrieval performance.
+As shown in Table 2, for the average performance across all magnifications, the employed medical foundation models deliver superior performances compared with all other CNN models, including DenseNet121.
+Based on the average results for the BreakHis dataset, the top-performing foundation model achieves a 1.7% and 4.5% improvement over the best CNN model in terms of mAP(5) and mMV(5), respectively.
+The results in Table 3 also indicate the superior performance of the medical foundation model feature extractors compared with other CNN models for the BACH dataset.
+Based on the results for the BACH dataset, the top-performing foundation model achieves a 7.7% and 15.5% improvement over the best CNN model in terms of mAP(5) and mMV(5), respectively.
+Overall, the results presented in Table 2 and Table 3 for the BreakHis and BACH datasets indicate that feature extractors trained on medical/histological images outperform those trained on non-medical images.
+Additionally, comparing foundation models shows highly competitive performance across all foundation models on the BreakHis dataset, while UNI achieves the best performance on the BACH dataset.
+In the second experiment, we compare the performance of the employed A-ARVGAE model with other GNN-based models for both BreakHis and BACH datasets.
 The results of these experiments are presented in Table 4.
-Table 4: Comparison of different models (GNN-based models including GAE, VGAE, and ARVGA, non-GNN-based model, and ours) for the BreakHis and BACH datasets for the breast cancer histological image retrieval task.
-GAE: graph autoencoder, VGAE: variational graph autoencoder, ARVGA: adversarial regularized variational graph autoencoder, RetCCL: retrieval with cluster-guided contrastive learning
+The UNI model was used as the graph feature extractor to derive the results in this table.
+Table 4: Comparison of different GNN-based models for the BreakHis and BACH datasets for the breast cancer histological image retrieval task.
+The UNI model was used as the graph feature extractor in all experiments.
+GAE: graph autoencoder, VGAE: variational graph autoencoder, ARVGA: adversarial regularized variational graph autoencoder, A-ARVGAE: attention-based adversarially regularized variational graph autoencoder.
 Dataset Method mAP(5) mMV(5)
-BreakHis (40$\times$) GAE 0.901 0.773
-VGAE 0.962 0.934
-ARVGA 0.969 0.909
-RetCCL 0.983 0.9789
-Ours 0.978 0.9499
-BreakHis (100$\times$) GAE 0.910 0.792
-VGAE 0.966 0.903
-ARVGA 0.949 0.909
-RetCCL 0.970 0.951
-Ours 0.980 0.961
-BreakHis (200$\times$) GAE 0.926 0.825
-VGAE 0.955 0.900
-ARVGA 0.949 0.879
-RetCCL 0.920 0.920
-Ours 0.964 0.930
-BreakHis (400$\times$) GAE 0.924 0.828
-VGAE 0.923 0.806
-ARVGA 0.932 0.834
-RetCCL 0.897 0.889
-Ours 0.940 0.839
-BreakHis Avg. GAE 0.915 0.804
-VGAE 0.951 0.885
-ARVGA 0.949 0.882
-RetCCL 0.942 0.934
-Ours 0.965 0.919
-BACH (200$\times$) GAE 0.919 0.740
-VGAE 0.937 0.798
-ARVGA 0.935 0.817
-RetCCL 0.901 0.8942
-Ours 0.947 0.913
-As shown in Table 4, the A-ARVGAE model surpasses all GNN-based models i.e., GAE, VGAE, and ARVGA for both datasets.
+BreakHis (40$\times$) GAE 0.897 0.783
+VGAE 0.960 0.894
+ARVGA 0.930 0.817
+A-ARVGAE 0.985 0.954
+BreakHis (100$\times$) GAE 0.896 0.768
+VGAE 0.960 0.937
+ARVGA 0.949 0.922
+A-ARVGAE 0.973 0.946
+BreakHis (200$\times$) GAE 0.907 0.771
+VGAE 0.930 0.855
+ARVGA 0.938 0.865
+A-ARVGAE 0.975 0.910
+BreakHis (400$\times$) GAE 0.925 0.784
+VGAE 0.934 0.790
+ARVGA 0.930 0.817
+A-ARVGAE 0.938 0.850
+BreakHis Avg. GAE 0.906 0.776
+VGAE 0.946 0.869
+ARVGA 0.936 0.855
+A-ARVGAE 0.967 0.915
+BACH (200$\times$) GAE 0.941 0.884
+VGAE 0.962 0.865
+ARVGA 0.962 0.855
+A-ARVGAE 0.976 0.942
+As shown in Table 4, the A-ARVGAE model surpasses all GNN-based models, i.e., GAE, VGAE, and ARVGA of both datasets.
 The superior performance of the proposed model over ARVGA demonstrates the significance of the attention component in the employed architecture.
 An attention component allows the model to focus on specific parts of input data to calculate the model weights.
-Compared with the RetCCL model, our approach delivers superior performance for both evaluation indices on the BACH dataset.
-However, for the BreakHis dataset, RetCLL shows slightly better performance for some magnifications.
-Nevertheless, our approach still yields superior results in other magnifications.
-The average results across all magnifications indicate the superior performance of our approach compared with RetCCL, based on the mAP(5) evaluation index (2.3% improved performance).
-However, for the MVP(5) score, RetCCL exhibits a better average performance with a 1.5% margin.
-Overall, while our approach has demonstrated superior or very competitive performance in all experiments, it should be noted that GNN training typically requires large datasets [59].
-The utilized model in this study was trained on the BreakHis and BACH datasets, which are relatively small.
+The average results indicate that the A-ARVGAE model achieves 2.1%/4.6% and 1.4%/5.8% higher performance in terms of mAP and mMV compared to the second-best GNN model for the BreakHis and BACH datasets.
 For visual analysis, we randomly select five query images from the BreakHis and BACH test sets (four from the BreakHis with different magnifications and one from the BACH dataset) to visualize the model output (top five ranks).
-Figure 5 depicts the query images and model output.
-As the results show, the model is capable of providing relevant images (with identical classes as the query images) for most of the samples.
+Fig.
+5 depicts the query images and model output, where the model (A-ARVGAE with UNI feature extractor) provides relevant images (with identical classes acting as the query images) for most samples.
 Query Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-M M M M M M
+BN M BN BN BN BN
 BH (100$\times$)
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-M BN M M M M
+BN BN M BN BN BN
 BH (200$\times$)
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-M M M M BN BN
+BN BN BN BN BN BN
 BH (400$\times$)
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-M M M M M M
+BN BN BN BN BN BN
 BACH (200$\times$)
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-N N N N N N
+BN BN BN BN BN BN
 Figure 5: Five sample queries from the BreakHis and BACH datasets and their top similar retrieved images.
-Successful and unsuccessful retrieval are denoted by () and (), respectively.
-M: Malignant class, BN: Benign class, N: Normal class, BH: BreakHis
-As suggested in [60], we also compare the embedding vectors of some example image pairs with different degrees of similarity in Figure 6.
-As shown in the figure, the embedding vectors for similar image pairs are very close to each other.
+Successful and unsuccessful retrieval are denoted by the symbols () and (), respectively.
+M: Malignant class, BN: Benign class, BH: BreakHis
+As suggested in [55], we also compare the embedding vectors of some example image pairs with different degrees of similarity in Fig.
+6, in which the embedding vectors for similar image pairs are very close.
 In contrast, the embedding vectors for dissimilar pairs are separable.
+Similar Pairs Refer to caption
+Refer to caption Refer to caption
+Dissimilar Pairs Refer to caption
+Refer to caption Refer to caption
+Similar Pairs Refer to caption
+Refer to caption Refer to caption
+Dissimilar Pairs Refer to caption
+Refer to caption Refer to caption
 Figure 6: Comparison of embedding values for four image pairs.
@@
 This research introduces a new GNN-based approach specifically designed for breast histopathology image retrieval.
-We incorporated both attention mechanisms and adversarial training in a variational graph auto-encoder model to enhance retrieval performance.
-Moreover, instead of using conventional pre-trained models as the node feature extractor, we employed the CCL feature extractor, specifically pre-trained on histological images.
-Our results confirm that incorporating both attention mechanisms and adversarial training consistently exhibits superior performance compared to other GNN variants.
-Additionally, using the CCL feature extractor further boosts performance in our experiments.
+We incorporated attention mechanisms and adversarial training in a variational graph auto-encoder model to enhance retrieval performance.
+Moreover, instead of using conventional pre-trained CNN models as the node feature extractor, we employed the features from medical foundation models, specifically pre-trained on medical/histological images.
+Our results confirm incorporating attention mechanisms and adversarial training outperforms the other GNN variants.
+Additionally, using the feature extracted from foundation models further boosts performance in our experiments.
 While our approach delivers promising results for breast histology image retrieval, certain aspects can be considered for further investigation in future studies.
@@
 Some Transformer-based GNNs, such as graph transformer networks [56], have been developed for tasks such as node classification.
-Such models can be adapted and employed for the medical image retrieval task as well.
-While we use FLANN to build the graph in our model, other state-of-the-art ANN methods and libraries, such as Faiss library  [62] or HNSW method  [8] can be exploited to enhance performance.
-However, it should be noted that most of these methods are slower in comparison to FLANN  [36].
+Such models can be adapted and employed for medical image retrieval.
+While we use FLANN to build the graph in our model, other state-of-the-art ANN methods and libraries, such as the Faiss library [57] or the HNSW method [8], can be exploited to enhance performance.
+However, most of these methods are slower than FLANN  [33].
 Another potential enhancement for the model is replacing the inner product decoder with a non-linear neural network, a topic that can be addressed in future studies.
-Finally, using well-known boosting techniques, such as test-time augmentation or ensembling, has been shown to enhance performance for tasks such as medical image classification or segmentation [63, 64, 65].
-Such techniques can also be incorporated into the workflow, and their impact on the retrieval performance can be investigated.
+We used three medical foundation models in the study as the graph node feature extractors.
+However, alternative foundation models could also be explored by considering recent advancements in this field and the availability of other medical and non-medical foundation models.
+We selected these specific medical foundation models because they have demonstrated superior performance than several other foundation models across various computer vision tasks in former studies  [21, 23, 58].
+Finally, well-known boosting techniques, such as test-time augmentation or ensembling, can enhance performance for medical image classification or segmentation  [59, 60, 61].
+The resulting retrieval performance can be investigated by incorporating such techniques into the workflow.
 Acknowledgments
@@
 Declaration of generative AI and AI-assisted technologies in the writing process
-During the preparation of this work, the authors used Grammarly (version 1.2.75) and ChatGPT (version 4) in order to check grammar and spelling and, at times, improve the readability of some sentences.
+During the preparation of this work, the authors used Grammarly (version 1.2.75) and ChatGPT (version 4) to check grammar and spelling and, at times, improve the readability of some sentences.
 After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.